OpenStudy (anonymous):

There were 1500 people at a high school football game. Student tickets were $2.00 and adult tickets were $3.50. The total receipts for the game were $3825. How many students bought tickets? • Let a = the number of adult tickets purchased. • Let s = the number of student tickets purchased. a. Write a system of equations that can be used to determine the number of adult and student

OpenStudy (andrewhaze):

.�

OpenStudy (anonymous):

let a be the number of adult ticket and s the number of children: We have the first equation is: a +s = 1500 (the number of adult plus children tickets) And for the second equation, we have: (3.50)a + (2)c=3825 From this point, there are several ways to solve this system, but it this case, let's do the elimination method (aka addition method) a+s=1500 3.50a+ 2s =3825 (We'll need to find the number that when multiply that number to either the first or second equation, one of the equation will cancel out and in this case, (-2)) so far, we have: (a+s=1500) x (-2) 3.50a +2s=3825 -2a -2s= -3000 3.50a+2s=3825 Now, add the two equations together, we'll have 1.5a= 825 ---->Solve for a, we'll have a=550----> Adult:550 But we're not quite done yet, we have to go back to either one of the equation and plug in the value for a (which is 550) and solve for the number of students a+s=1500 (550)+s=1500 s=950 The answer is, I believe, 950 students bought the tickets.